OpenStudy (anonymous):

you accidentally dropped a coin from the top of 12 stairs. what is the probability that it will land on the tenth step,heads up?

OpenStudy (anonymous):



OpenStudy (dumbcow):

hmm , doesn't matter where you drop a coin, the probability it lands heads up is 1/2

OpenStudy (anonymous):

what does it says on it?

OpenStudy (anonymous):

is dat 5/12

OpenStudy (dumbcow):

oh i see now...yes that is 5/12 and it is correct

OpenStudy (anonymous):

r u sure??

OpenStudy (dumbcow):

prob coin lands on 10th step = 10/12 prob lands heads = 1/2 multiply 2 independent event: (10/12)(1/2) = 5/12